Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

four digit safe code

Expert replies
by gmataug08 » Mon Nov 24, 2008 7:45 am
A Four digit safe code does not contain the digits 1 and 4 at all. What is the probability that it has at least one even digit?

a) ¼
b) ½
c) ¾
d) 15/16
e) 1/16

OA : D
Join the discussion
Source: — Problem Solving |

by mals24 » Mon Nov 24, 2008 8:06 am
Probability that code has at least 1 even digit = 1 - Prob. code does not have any even digit.

Numbers available = 0,2,3,5,6,7,8,9
Total number of odd numbers = 4 (3,5,7,9)

Now since the question does not mention that all the digits in the code are different, we can repeat the numbers.

There are 4 options for the 1st digit.
4 options for the 2nd
4 for the 3rd
4 for the fourth.

Number of ways of selecting odd digits = 4*4*4*4
Number of ways of selecting 4 digits = 8*8*8*8

Total probability that all digits are odd = 4^4/8^4 = 1/16

Hence prob. that at least 1 digit is even = 1-1/16 = 15/16
Join the discussion

by moadhia » Mon Nov 24, 2008 8:06 am
The 4 digit code looks like this

_ _ _ _


The question is to find the probability of atleast one even digit, the easiest way to do this would be to find the probability of having no even digit and subtracting it by 1

each digit can have 8 different values so denominator = 8 * 8 * 8 *8 (1)

for all the digits to have only odd numbers the # of ways = 4 * 4 * 4 * 4 (2)

prob of having no even digit = 1 / 16 ( Divide (2) by (1) )


1 - 1/16 = 15/16

Hope this helps.
Join the discussion

Re: four digit safe code

by Stuart@KaplanGMAT » Mon Nov 24, 2008 10:35 am
gmataug08 wrote:A Four digit safe code does not contain the digits 1 and 4 at all. What is the probability that it has at least one even digit?

a) ¼
b) ½
c) ¾
d) 15/16
e) 1/16
Understanding probability should make this a 15 second question with a 99% confidence level.

Let's use some common sense. The 8 remaining digits are evenly distributed between even/odd. Right away we know that there will be more than a 50% chance of at least one even digit: eliminate a, b and e.

We know there are 4 digits, so the chance of getting at least one even should be extremely high: out of c and d, d is clearly a better choice.

If you want to actually solve, the solutions posted above are both correct. We could also recognize that since there's a 50% chance that each digit is even (4 evens, 4 odds), this is really a coin flip question, as follows:
If you flip a fair coin 4 times, what's the probability of getting at least 1 head?
Since the only results we do NOT want is 0 heads, we use the "1 minus" approach:

Prob(want) = 1 - Prob(don't want)

Here's the generic coin flip formula:

Prob(k results from n flips) = nCk/2^n

With k=0 and n=4, we get:

4C0/2^4 = 1/16

(some good combinations shortcuts to memorize:
nC0 = 1
nC1 = n
nCn = 1)

So:

Prob(at least 1H) = 1 - 1/16 = 15/16
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion